OpenStudy (anonymous):

Evaluate the integral using integration by parts where possible: (x)/((x-9)^2)dx

OpenStudy (anonymous):

why parts? partial fraction or u-sub will get it right away

OpenStudy (anonymous):

hey satellite i was wondering if you could help me after?

OpenStudy (anonymous):

let me try that again \[u=x+9, du = dx, x=u+9\] \[\int\frac{u+9}{u^2}du=\int \frac{1}{u}du+\int\frac{9}{u^2}du\] \[\ln(u)-\frac{9}{u}\] \[\ln(x-9)-\frac{9}{(x-9)}\]